# AdapDISCOM: An Adaptive Sparse Regression Method for High-Dimensional Multimodal Data With Block-Wise Missingness and Measurement Errors

### v1

@@
Abdoul O. Diakité Institut national de la recherche scientifique (INRS), Centre Armand-Frappier Santé Biotechnologie, Laval, Canada Unité mixte de recherche (UMR) INRS-UQAC en santé durable, Chicoutimi, Canada Claudia Moreau Unité mixte de recherche (UMR) INRS-UQAC en santé durable, Chicoutimi, Canada Université du Québec à Chicoutimi (UQAC), Chicoutimi, Canada Gleb Bezgin Department of Neurology and Neurosurgery, Faculty of Medicine and Health Sciences, McGill University, Montreal, Quebec, Canada Nikhil Bhagwat NeuroDataScience-ORIGAMI lab, McConnell Brain Imaging Centre, The Neuro (Montreal Neurological Institute-Hospital), Faculty of Medicine and Health Sciences, McGill University, Montreal, Quebec, Canada Pedro Rosa-Neto Translational Neuroimaging Laboratory, McGill University Research Centre for Studies in Aging, McConnell Brain Imaging Centre (BIC), Montreal Neurological Institute, Montreal Neurological Institute-Hospital, Montreal, QC, Canada Institution: Douglas Hospital Research Centre - Centre intégré universitaire de santé et services sociaux de l’Ouest-de-l’Île-de-Montréal, Verdun, Quebec, Canada The Peter O’Donnell Jr. Brain Institute (OBI), University of Texas Southwestern Medical Centre (UTSW). Dallas, TX, USA Jean-Baptiste Poline NeuroDataScience-ORIGAMI lab, McConnell Brain Imaging Centre, The Neuro (Montreal Neurological Institute-Hospital), Faculty of Medicine and Health Sciences, McGill University, Montreal, Quebec, Canada Simon Girard Unité mixte de recherche (UMR) INRS-UQAC en santé durable, Chicoutimi, Canada Université du Québec à Chicoutimi (UQAC), Chicoutimi, Canada Amadou Barry [^1] Corresponding author: amadoudiogo.barry@inrsc.ca. Institut national de la recherche scientifique (INRS), Centre Armand-Frappier Santé Biotechnologie, Laval, Canada Unité mixte de recherche (UMR) INRS-UQAC en santé durable, Chicoutimi, Canada for the Alzheimer’s Disease Neuroimaging Initiative Data used in preparation of this article were obtained from the Alzheimer’s Disease Neuroimaging Initiative (ADNI) database (adni.loni.usc.edu). As such, the investigators within the ADNI contributed to the design and implementation of ADNI and/or provided data but did not participate in analysis or writing of this report. A complete listing of ADNI investigators can be found at: http://adni.loni.usc.edu/wp-content/uploads/how_to_apply/ADNI_Acknowledgement_List.pdf
(July 31, 2025)
@@ 2 Introduction
Multimodal, high-dimensional data are increasingly pervasive in modern scientific research, underpinning advances in fields such as health, genetics, and neuroscience. By integrating complementary information from diverse modalities, such datasets enable richer modeling and more accurate prediction, supporting personalized medicine, large-scale genetic studies, and drug discovery [19]. However, their analysis presents significant statistical and computational challenges, particularly when data are subject to block-wise missingness and measurement errors —two common imperfections in real-world settings.
@@
In biobanks like the Alzheimer’s Disease Neuroimaging Initiative (ADNI) study [35], which integrates clinical, genetic, and neuroimaging data (structural, functional, molecular), block-wise missingness arises when entire modalities are unavailable for some individuals—often due to high costs (e.g., PET scans), technical failures, or patient reluctance (e.g., CSF collection). Simultaneously, the observed data are frequently contaminated with additive measurement errors, whose magnitude and nature vary across modalities depending on acquisition techniques. Ignoring these issues risks producing biased estimates, poor predictive performance, and unreliable variable selection. Standard statistical methods and algorithms fail to address these challenges effectively. Therefore, developing innovative methods and algorithms to mitigate the influence of measurement errors and missing data is crucial, especially given the size and complexity of modern datasets.
@@
In recent years, substantial research efforts have been devoted to reducing the impact of measurement error and missing values in high-dimensional data. Regarding measurement errors, Loh and Wainwright, [29] introduced a nonconvex modification of the Lasso objective function [45], incorporating a constrained parameter that adds an optimization constraint. Despite providing theoretical guarantees of convergence close to the global minimum, the nonconvex nature of the objective function complicates computations. To address this limitation, Datta and Zou, [8] proposed the Convex Conditioned Lasso (CoCoLasso), utilizing a positive semi-definite covariance estimator to achieve similar theoretical and computational advantages as the Lasso. Subsequent improvements have been proposed, such as BD-CoCoLasso [11], which reduces the computational burden, and balanced estimation techniques [58], which introduce an $l_{0}$ penalty to CoCoLasso for improved variable selection and prediction trade-offs. More recently, Tao et al., [44] proposed the CaZnRLS estimator, which calibrates least squares loss with a positively defined projection of an unbiased substitute for the covariate covariance matrix. A comprehensive review of recent advances in measurement error correction methods for high-dimensional data can be found in [30].
@@
Notably, CoCoLasso and its extensions primarily provide separate solutions for multiplicative measurement errors, where missing data are treated as a particular case. Meanwhile, recent methodologies have focused explicitly on handling missing data in high-dimensional settings. Many of these methods depart from traditional imputation approaches, which generate unstable matrices, accumulate imputation errors, and are computationally expensive for large-scale data. For instance, Li et al., [28] proposed a sample-wise combined missing effect model (SCOM) with penalization, integrating missing data from each sample into the model specification as an estimable parameter. Yu et al., [54] developed a direct sparse regression procedure using covariance from multimodality (DISCOM) method, inspired by Ledoit and Wolf, [27]’s covariance estimator, to account for missing data in multimodal high-dimensional datasets. DISCOM estimates covariance as a linear combination of the identity matrix, intra-modal covariance estimates, and cross-modal covariance estimates. This covariance estimator is then incorporated into Lasso optimization for sparse estimation of optimal linear prediction coefficients. A key advantage of DISCOM is its ability to leverage all available information, as it only requires the minimum number of samples with at least two observed modalities, significantly increasing the effective sample size compared to complete-case analysis. DISCOM has gained considerable interest, with recent adaptations for multi-omics network data [23], accelerated failure time models with censored responses [51], and multi-response variable settings [52]. Additionally, He et al., [22] integrated DISCOM into an imputation algorithm for generalized linear models, further extending its applicability.
@@ 3.1 DISCOM accounting for missing data and additive measurement error
We want to predict the centered response variable y from the linear regression model $\textbf{y}=\textbf{X}\boldsymbol{\beta}+\boldsymbol{\varepsilon},$ where $\textbf{X}=\big{[}\textbf{X}^{(1)},\textbf{X}^{(2)},\ldots,\textbf{X}^{(K)}\big{]}$ is the $K$ multimodal matrix of predictors, $\boldsymbol{\beta}\in\mathbb{R}^{p}$ the associated parameter, and $\boldsymbol{\varepsilon}$ the random error. For this purpose, we consider a data sample $\{\textbf{y}=(y_{i}),\textbf{X}=(x_{ij})\}_{1\leq i\leq n,1\leq j\leq p}$ of size $n$ and $p$ predictors generated from some multivariate distribution with mean $\boldsymbol{0}_{p\times 1}$ and covariance matrix $\Sigma.$ Note that, each of the $K$ modalities have $p_{k}$ predictors for $k\in\{1,2,\ldots,K\}.$ As in high-dimensional settings, where sparsity assumptions are commonly made [21, 5], we assume that only a finite set of predictors contributes to the prediction of the response variable. Accordingly, to obtain a sparse estimator of the true regression coefficient $\boldsymbol{\beta}^{0}\in\mathbb{R}^{p},$ we solve the following penalized optimization problem:
@@
[…]
@@
Without loss of generality, we adopt the framework of [54] and assume a block-missing multimodality data. For each sample, if a certain modality has missing entries, all the observations from that modality are missing. Assuming a missing completely at random mechanism for each modality, we consider as a natural initial estimate of $\Sigma$ the covariance matrix $\widetilde{\Sigma}=\frac{1}{n}\textbf{Z}{}^{\text{T}}\textbf{Z}=(\widetilde{\sigma}_{jt})_{j,t=1,\ldots,p}$ where $\widetilde{\sigma}_{jt}=\frac{1}{n_{jt}}\sum_{i\in\text{S}_{jt}}z_{ij}z_{it}$ is estimated using all the available data. The set $\text{S}_{jt}=\{i:z_{ij}\ \text{and}\ z_{it}\ \text{are not missing}\}$ represent the set of individuals with non missing data for the pair of predictors $j$ and $t$ and $n_{jt}$ is the cardinal number of $\text{S}_{jt}.$ When $j=t$ then $\text{S}_{jt}=\text{S}_{j}$ and $n_{jt}=n_{j}.$ However, $\widetilde{\Sigma}$ may have negative eigenvalues due to the unequal sample sizes and is not a good estimate of the covariance matrix $\Sigma,$ and not suitable to be used in (2) directly [54].
@@
Following the observation of [54], we note that the covariance matrix $\widetilde{\Sigma}$ can be partitioned into the form $\widetilde{\Sigma}=\widetilde{\Sigma}_{I}+\widetilde{\Sigma}_{C}$ where $\widetilde{\Sigma}_{I}=\operatorname{diag}(\widetilde{\Sigma}_{11},\ \widetilde{\Sigma}_{22},\ldots,\widetilde{\Sigma}_{KK})=\operatorname{diag}(\widetilde{\Sigma}_{I1},\ \widetilde{\Sigma}_{I2},\ldots,\widetilde{\Sigma}_{IK})$ is the intra-modality sample covariance matrix, and $\widetilde{\Sigma}_{C}=\widetilde{\Sigma}-\widetilde{\Sigma}_{I}$ the cross-modality sample covariance matrix. The matrix $\widetilde{\Sigma}_{C}$ is an off-diagonal matrix and contains $K(K-1)$ blocks matrix denoted by $\widetilde{\Sigma}_{jt}$’s, where $j,\ t\in\{1,2,\ldots,p\}$ and $\widetilde{\Sigma}_{jt}$ is a $p_{j}\times p_{t}$ matrix. We also note $\Sigma_{I}=\operatorname{diag}(\Sigma_{I1},\ \Sigma_{I2},\ldots,\Sigma_{IK})$ and $\Sigma_{C}$ the true intra-modality and cross-modality covariance matrix. Likewise, $\Sigma_{jt}$ is the true covariance matrix of $\widetilde{\Sigma}_{jt}$ for $j,\ t\in\{1,2,\ldots,p\}.$
@@
We now address the mitigation of measurement error and missing data issues in two sequential steps, starting with the case of additive measurement error. In the presence of additive measurement error, Datta and Zou, [8] proposed the CocoLasso estimator $\grave{\Sigma}=\widetilde{\Sigma}+\gamma^{2}\mathbb{I}_{p}$ as an unbiased estimator of the covariance matrix, where $\gamma^{2}$ is the variance of the additive error matrix [29]. By applying the CocoLasso estimator separately to the intra-modality covariance matrix $\grave{\Sigma}_{I}=\widetilde{\Sigma}_{I}+\gamma^{2}_{I}\mathbb{I}_{p}$ and cross-modality covariance matrix $\grave{\Sigma}_{C}=\widetilde{\Sigma}_{C}+\gamma^{2}_{C}\mathbb{I}_{p},$ we can mitigate the additive measurement errors and then solve the following optimization problem
@@
However, as the CoCoLasso covariance estimator only mitigates the influence of the additive measurement error, the estimator derived from (3) will still be biased due to the presence of missing data. As a second step, we combine the DISCOM estimator [54] to the CocoLasso covariance estimator in order to mitigate the presence of missing data and ultimately reduce simultaneously the influence of both issues. The DISCOM covariance estimator for multimodal data is a linear combination of the intra-modality covariance matrix, the cross-modality covariance matrix and the identity matrix. Thus, the covariance estimator that takes into account both measurement errors and missing data is defined by

### v2

@@
Maimouna Baldé Institut national de la recherche scientifique (INRS), Centre Armand-Frappier Santé Biotechnologie, Laval, Canada Unité mixte de recherche (UMR) INRS-UQAC en santé durable, Chicoutimi, Canada Abdoul O. Diakité Institut national de la recherche scientifique (INRS), Centre Armand-Frappier Santé Biotechnologie, Laval, Canada Unité mixte de recherche (UMR) INRS-UQAC en santé durable, Chicoutimi, Canada Claudia Moreau Unité mixte de recherche (UMR) INRS-UQAC en santé durable, Chicoutimi, Canada Université du Québec à Chicoutimi (UQAC), Chicoutimi, Canada Gleb Bezgin Department of Neurology and Neurosurgery, Faculty of Medicine and Health Sciences, McGill University, Montreal, Quebec, Canada Nikhil Bhagwat NeuroDataScience-ORIGAMI lab, McConnell Brain Imaging Centre, The Neuro (Montreal Neurological Institute-Hospital), Faculty of Medicine and Health Sciences, McGill University, Montreal, Quebec, Canada Pedro Rosa-Neto Translational Neuroimaging Laboratory, McGill University Research Centre for Studies in Aging, McConnell Brain Imaging Centre (BIC), Montreal Neurological Institute, Montreal Neurological Institute-Hospital, Montreal, QC, Canada Institution: Douglas Hospital Research Centre - Centre intégré universitaire de santé et services sociaux de l’Ouest-de-l’Île-de-Montréal, Verdun, Quebec, Canada The Peter O’Donnell Jr. Brain Institute (OBI), University of Texas Southwestern Medical Centre (UTSW). Dallas, TX, USA Jean-Baptiste Poline NeuroDataScience-ORIGAMI lab, McConnell Brain Imaging Centre, The Neuro (Montreal Neurological Institute-Hospital), Faculty of Medicine and Health Sciences, McGill University, Montreal, Quebec, Canada Simon Girard Unité mixte de recherche (UMR) INRS-UQAC en santé durable, Chicoutimi, Canada Université du Québec à Chicoutimi (UQAC), Chicoutimi, Canada Amadou Barry [^1] Corresponding author: amadoudiogo.barry@inrsc.ca. Institut national de la recherche scientifique (INRS), Centre Armand-Frappier Santé Biotechnologie, Laval, Canada Unité mixte de recherche (UMR) INRS-UQAC en santé durable, Chicoutimi, Canada for the Alzheimer’s Disease Neuroimaging Initiative Data used in preparation of this article were obtained from the Alzheimer’s Disease Neuroimaging Initiative (ADNI) database (adni.loni.usc.edu). As such, the investigators within the ADNI contributed to the design and implementation of ADNI and/or provided data but did not participate in analysis or writing of this report. A complete listing of ADNI investigators can be found at: http://adni.loni.usc.edu/wp-content/uploads/how_to_apply/ADNI_Acknowledgement_List.pdf
(February 4, 2026)
@@ 2 Introduction
Multimodal, high-dimensional data are increasingly pervasive in modern scientific research, underpinning advances in fields such as health, genetics, and neuroscience. By integrating complementary information from diverse modalities, such datasets enable richer modeling and more accurate prediction, supporting personalized medicine, large-scale genetic studies, and drug discovery [García and Purriños, 2014]. However, their analysis presents significant statistical and computational challenges, particularly when data are subject to block-wise missingness and measurement errors —two common imperfections in real-world settings.
@@
In biobanks like the Alzheimer’s Disease Neuroimaging Initiative (ADNI) study [Mueller et al., 2005b], which integrates clinical, genetic, and neuroimaging data (structural, functional, molecular), block-wise missingness arises when entire modalities are unavailable for some individuals—often due to high costs (e.g., PET scans), technical failures, or patient reluctance (e.g., CSF collection). Simultaneously, the observed data are frequently contaminated with additive measurement errors, whose magnitude and nature vary across modalities depending on acquisition techniques. Ignoring these issues risks producing biased estimates, poor predictive performance, and unreliable variable selection. Standard statistical methods and algorithms fail to address these challenges effectively. Therefore, developing innovative methods and algorithms to mitigate the influence of measurement errors and missing data is crucial, especially given the size and complexity of modern datasets.
@@
In recent years, substantial research efforts have been devoted to reducing the impact of measurement error and missing values in high-dimensional data. Regarding measurement errors, Loh and Wainwright [2012] introduced a nonconvex modification of the Lasso objective function [Tibshirani, 1996], incorporating a constrained parameter that adds an optimization constraint. Despite providing theoretical guarantees of convergence close to the global minimum, the nonconvex nature of the objective function complicates computations. To address this limitation, Datta and Zou [2017] proposed the Convex Conditioned Lasso (CoCoLasso), utilizing a positive semi-definite covariance estimator to achieve similar theoretical and computational advantages as the Lasso. Subsequent improvements have been proposed, such as BD-CoCoLasso [Escribe et al., 2021], which reduces the computational burden, and balanced estimation techniques [Zheng et al., 2018], which introduce an $l_{0}$ penalty to CoCoLasso for improved variable selection and prediction trade-offs. More recently, Tao et al. [2021] proposed the CaZnRLS estimator, which calibrates least squares loss with a positively defined projection of an unbiased substitute for the covariate covariance matrix. A comprehensive review of recent advances in measurement error correction methods for high-dimensional data can be found in [Luo et al., 2023].
@@
Notably, CoCoLasso and its extensions primarily provide separate solutions for multiplicative measurement errors, where missing data are treated as a particular case. Meanwhile, recent methodologies have focused explicitly on handling missing data in high-dimensional settings. Many of these methods depart from traditional imputation approaches, which generate unstable matrices, accumulate imputation errors, and are computationally expensive for large-scale data. For instance, Li et al. [2023] proposed a sample-wise combined missing effect model (SCOM) with penalization, integrating missing data from each sample into the model specification as an estimable parameter. Yu et al. [2020] developed a direct sparse regression procedure using covariance from multimodality (DISCOM) method, inspired by Ledoit and Wolf [2004]’s covariance estimator, to account for missing data in multimodal high-dimensional datasets. DISCOM estimates covariance as a linear combination of the identity matrix, intra-modal covariance estimates, and cross-modal covariance estimates. This covariance estimator is then incorporated into Lasso optimization for sparse estimation of optimal linear prediction coefficients. A key advantage of DISCOM is its ability to leverage all available information, as it only requires the minimum number of samples with at least two observed modalities, significantly increasing the effective sample size compared to complete-case analysis. DISCOM has gained considerable interest, with recent adaptations for multi-omics network data [Henao et al., 2023], accelerated failure time models with censored responses [Wang et al., 2022], and multi-response variable settings [Wang et al., 2024]. Additionally, He et al. [2023] integrated DISCOM into an imputation algorithm for generalized linear models, further extending its applicability.
@@ 3.1 DISCOM accounting for missing data and additive measurement error
We want to predict the centered response variable y from the linear regression model $\textbf{y}=\textbf{X}\boldsymbol{\beta}+\boldsymbol{\varepsilon},$ where $\textbf{X}=\big[\textbf{X}^{(1)},\textbf{X}^{(2)},\ldots,\textbf{X}^{(K)}\big]$ is the $K$ multimodal matrix of predictors, $\boldsymbol{\beta}\in\mathbb{R}^{p}$ the associated parameter, and $\boldsymbol{\varepsilon}$ the random error. For this purpose, we consider a data sample $\{\textbf{y}=(y_{i}),\textbf{X}=(x_{ij})\}_{1\leq i\leq n,1\leq j\leq p}$ of size $n$ and $p$ predictors generated from some multivariate distribution with mean $\boldsymbol{0}_{p\times 1}$ and covariance matrix $\Sigma.$ Note that, each of the $K$ modalities have $p_{k}$ predictors for $k\in\{1,2,\ldots,K\}.$ As in high-dimensional settings, where sparsity assumptions are commonly made [Hastie et al., 2009; Bühlmann and van de Geer, 2011], we assume that only a finite set of predictors contributes to the prediction of the response variable. Accordingly, to obtain a sparse estimator of the true regression coefficient $\boldsymbol{\beta}^{0}\in\mathbb{R}^{p},$ we solve the following penalized optimization problem:
@@
where $\widehat{\boldsymbol{\beta}}$ is the estimator of the true unknown parameter $\boldsymbol{\beta}^{0},\ \lambda>0$ a tuning parameter and $\mathcal{L}(\cdot)$ a generic penalty function such as LASSO [Tibshirani, 1996], used to obtain a sparse solution. Note that, optimization of equation (1) is equivalent to solving: $\operatornamewithlimits{argmin}_{\boldsymbol{\beta}\in\mathbb{R}^{p}}\frac{1}{2}\boldsymbol{\beta}{}^{\text{T}}\Sigma\boldsymbol{\beta}+\boldsymbol{\beta}{}^{\text{T}}\textbf{C}+\mathcal{L}(\boldsymbol{\beta}),$ where $\Sigma=\frac{1}{n}\textbf{X}{}^{\text{T}}\textbf{X}$ is the positive semi-definite covariance matrix and $\textbf{C}=\frac{1}{n}\textbf{X}{}^{\text{T}}\textbf{y}$ the covariance vector between the predictors and the response variable. However, in practice, the matrix $\textbf{X}=(x_{ij})_{1\leq i\leq n,1\leq j\leq p}$ is not directly observable, especially in high dimensions, and consequently $\Sigma$ and $\textbf{C}.$ Instead, we have a multimodal matrix $\textbf{Z}=\big[\textbf{Z}^{(1)},\textbf{Z}^{(2)},\ldots,\textbf{Z}^{(K)}\big]=(z_{ij})_{1\leq i\leq n,1\leq j\leq p},$ affected simultaneously by additive measurement errors and missing data, and we need to find an unbiased estimates $\widehat{\Sigma}$ of $\Sigma$ and $\widehat{\textbf{C}}$ of C to solve the following problem:
@@
Without loss of generality, we adopt the framework of [Yu et al., 2020] and assume a block-missing multimodality data. For each sample, if a certain modality has missing entries, all the observations from that modality are missing. Assuming a missing completely at random mechanism for each modality, we consider as a natural initial estimate of $\Sigma$ the covariance matrix $\widetilde{\Sigma}=\frac{1}{n}\textbf{Z}{}^{\text{T}}\textbf{Z}=(\widetilde{\sigma}_{jt})_{j,t=1,\ldots,p}$ where $\widetilde{\sigma}_{jt}=\frac{1}{n_{jt}}\sum_{i\in\text{S}_{jt}}z_{ij}z_{it}$ is estimated using all the available data. The set $\text{S}_{jt}=\{i:z_{ij}\ \text{and}\ z_{it}\ \text{are not missing}\}$ represent the set of individuals with non missing data for the pair of predictors $j$ and $t$ and $n_{jt}$ is the cardinal number of $\text{S}_{jt}.$ When $j=t$ then $\text{S}_{jt}=\text{S}_{j}$ and $n_{jt}=n_{j}.$ However, $\widetilde{\Sigma}$ may have negative eigenvalues due to the unequal sample sizes and is not a good estimate of the covariance matrix $\Sigma,$ and not suitable to be used in (2) directly [Yu et al., 2020].
@@
Following the observation of [Yu et al., 2020], we note that the covariance matrix $\widetilde{\Sigma}$ can be partitioned into the form $\widetilde{\Sigma}=\widetilde{\Sigma}_{I}+\widetilde{\Sigma}_{C}$ where $\widetilde{\Sigma}_{I}=\operatorname{diag}(\widetilde{\Sigma}_{11},\ \widetilde{\Sigma}_{22},\ldots,\widetilde{\Sigma}_{KK})=\operatorname{diag}(\widetilde{\Sigma}_{I1},\ \widetilde{\Sigma}_{I2},\ldots,\widetilde{\Sigma}_{IK})$ is the intra-modality sample covariance matrix, and $\widetilde{\Sigma}_{C}=\widetilde{\Sigma}-\widetilde{\Sigma}_{I}$ the cross-modality sample covariance matrix. The matrix $\widetilde{\Sigma}_{C}$ is an off-diagonal matrix and contains K(K1)K(K-1) blocks matrix denoted by $\widetilde{\Sigma}_{jt}$’s, where $j,\ t\in\{1,2,\ldots,p\}$ and $\widetilde{\Sigma}_{jt}$ is a $p_{j}\times p_{t}$ matrix. We also note $\Sigma_{I}=\operatorname{diag}(\Sigma_{I1},\ \Sigma_{I2},\ldots,\Sigma_{IK})$ and $\Sigma_{C}$ the true intra-modality and cross-modality covariance matrix. Likewise, $\Sigma_{jt}$ is the true covariance matrix of $\widetilde{\Sigma}_{jt}$ for $j,\ t\in\{1,2,\ldots,p\}.$
@@
We now address the mitigation of measurement error and missing data issues in two sequential steps, starting with the case of additive measurement error. In the presence of additive measurement error, Datta and Zou [2017] proposed the CocoLasso estimator $\grave{\Sigma}=\widetilde{\Sigma}+\gamma^{2}\mathbb{I}_{p}$ as an unbiased estimator of the covariance matrix, where $\gamma^{2}$ is the variance of the additive error matrix [Loh and Wainwright, 2012]. By applying the CocoLasso estimator separately to the intra-modality covariance matrix $\grave{\Sigma}_{I}=\widetilde{\Sigma}_{I}+\gamma^{2}_{I}\mathbb{I}_{p}$ and cross-modality covariance matrix $\grave{\Sigma}_{C}=\widetilde{\Sigma}_{C}+\gamma^{2}_{C}\mathbb{I}_{p},$ we can mitigate the additive measurement errors and then solve the following optimization problem
@@
However, as the CoCoLasso covariance estimator only mitigates the influence of the additive measurement error, the estimator derived from (3) will still be biased due to the presence of missing data. As a second step, we combine the DISCOM estimator [Yu et al., 2020] to the CocoLasso covariance estimator in order to mitigate the presence of missing data and ultimately reduce simultaneously the influence of both issues. The DISCOM covariance estimator for multimodal data is a linear combination of the intra-modality covariance matrix, the cross-modality covariance matrix and the identity matrix. Thus, the covariance estimator that takes into account both measurement errors and missing data is defined by